# Self-Consistent Renormalized Spin-Wave Theory of Magnetic and Topological Transitions in Two-Dimensional Honeycomb Ferromagnets

### v1

@@ I Introduction
Since the discovery of intrinsic ferromagnetism in two-dimensional (2D) materials, this subject has attracted extensive attention due to its potential applications in novel spintronics devices [Gong2017, Huang17]. In particular, magnons in these 2D Van der Waals ferromagnetic or antiferromagnetic materials play a pivotal role in transport properties such as thermal Hall effects [Zhang24, Neumann22, Saleem24]. Moreover, since magnons are charge-neutral, they can propagate without experiencing any joule heating dissipation [Zhuo23]. Candidate materials include $\rm{Cr_{2}Ge_{2}Te_{6}}$ [Gong2017], $\rm{CrI_{3}}$ [Chen18, Chen21], $\rm{CrBr_{3}}$ [Zhang2019], $\rm{FePSe_{3}}$ [Luo23], etc. Recent theoretical investigations of topological properties in 2D magnonic systems [Pershoguba18, Habel24, Chen24, Hua23], particularly under the influence of magnon-magnon interaction within honeycomb ferromagnets, also mark an important advancement in the understanding and manipulation of topological phases of Dirac magnons [Lu2021, Mook2021, Mkhitaryan2021]. The influence of magnon-magnon interaction on honeycomb antiferromagnets has also been extensively investigated recently [Chen23, Li23, Neumann22, Sourounis24].
@@
In 2D topological magnetic materials, the crucial ingredient in achieving nontrivial topological properties is the Dzyaloshinskii-Moriya interaction (DMI) [Buzo24, Lu2021, Mook2021, Brehm24, Ni24, Zhu23, Kim22, Zhang23] which plays a role similar to the spin-orbit interaction in topological electronic systems [Haldane88]. Both Refs. [Lu2021] and [Mook2021] found the phenomenon of magnon-magnon interaction-induced topological phase transitions accompanied with sign reversals of the thermal Hall conductivity in honeycomb ferromagnets. These transitions can be induced by tuning the magnetic field or varying the temperature. Although the predicted phenomenon is similar in Refs. [Lu2021] and [Mook2021], the underlying setup is different in terms of the chosen directions of magnetic field and types of DMI. Alternatively, the topological phase transitions can also be controlled by chirality injection from boundaries [Lee23], manipulation of the ground-state magnetization direction [Klogetvedt23], or parametric magnon amplification scheme [Sun23], etc. In addition to honeycomb lattices, we note here that topology in 2D magnetic materials is also important in kagome lattices [Mook14, Hirschberger15, Seshardi18, Owerre18] and triangular kagome lattices [Zhang23].
@@
A common description of magnon behavior in ferromagnets primarily relies on spin-wave theory. When interactions between magnons are taken into account, intriguing phenomena such as topological phase transitions can emerge. The theory presented in Refs. [Lu2021, Mook2021] employs the standard Green’s function approach; however, Ref. [Lu2021] only considers the case where the proper self-energy of the Hartree type is composed of bare Green’s function. This means that magnon-magnon interactions are not treated in a fully self-consistent manner in Ref. [Lu2021]. It is suggested in Ref. [Li24] that, to properly compare the temperatures of topological phase transitions ($T_{c}$) with Curie temperatures ($T_{\rm Curie}$), the self-consistently renormalized spin-wave theory (SRSWT) should be used, as done in Ref. [Mkhitaryan2021]. Here, we refer to the Curie temperature as the point at which magnetic order disappears, regardless of whether the underlying transition is continuous or first-order, as may arise within SRSWT. The corresponding self-energy in SRSWT [Bloch62, Loly71, Rastelli74, Pini81, Li18] is equivalent to summing over all bubble diagrams and depends on the dressed Green’s function. Even though the SRSWT allows extending the highest temperature to the order of $T_{\rm Curie}$, the theory still has some inherent drawbacks. For example, a key weakness is the predicted first-order ferromagnetic phase transition, whereas it is known to be second-order [Rastelli74, Li18]. This drawback arises from the violation of the kinematical constraints [Irkhin97, Mkhitaryan2021] at high temperatures.
@@
To clarify the hierarchy of approximations, it is useful to contrast the different levels of spin-wave theory. In linear spin-wave theory (LSWT), magnon–magnon interactions are entirely neglected, leading to a systematic overestimate of $T_{\rm Curie}$ [Li18]. A strict $1/S$ expansion incorporates interaction terms, but evaluates self-energy diagrams using bare Green’s functions; this provides controlled corrections at intermediate temperatures. By contrast, SRSWT employs dressed propagators in the same diagrams, thereby extending the applicability of spin-wave theory up to $T_{\rm Curie}$ but at the expense of artifacts such as a spurious first-order–like transition, which do not arise in the strict $1/S$ expansion.
@@
If the SRSWT is adopted for the problem, it was argued in Ref. [Li24] that the temperature at which the magnetization vanishes coincides with the Curie temperature. As a result, topological phase transitions would be inaccessible in transport or inelastic neutron scattering experiments. In the corresponding Reply [Lu24], we incorporated the next-order contribution to the magnon self-energy arising from the Holstein–Primakoff (HP) expansion, and demonstrated that, even within the SRSWT framework, the topological transition occurs at a finite magnetization. Motivated by the findings in the Reply [Lu24], the present work aims to present the calculation methods in greater detail, offer a more comprehensive analysis of the conditions under which topological phase transitions can occur—specifically, at temperatures below $T_{\rm Curie}$—and critically examine the applicability of SRSWT to Dirac magnons.
@@ II Methods
Figure 1: (a) The illustration of a honeycomb lattice, showing the three nearest-neighbor and next-nearest-neighbor vectors, labeled as $\bm{\delta}_{i}$ and $\bm{\zeta}_{i}$ ($i=1,2,3$), respectively. We set the out-of-plane direction to be along $\hat{\bm{z}}$. The blue and red sites are denoted as AA and $B$ sites, respectively. (b) Bubble diagrams corresponding to Hartree-type self-energy contributions in the self-consistent renormalized spin-wave theory. (c) Feynman diagram representing six-operators interaction terms.
@@
The first two terms in the Hamiltonian represent the conventional Heisenberg interactions for nearest-neighbor and next-nearest-neighbor pairs. The third term corresponds to the out-of-plane NNN Dzyaloshinskii-Moriya interaction, denoted by $\bm{D}_{ij}=\pm D\hat{\bm{z}}$, where the plus (minus) sign corresponds to clockwise (counterclockwise) hopping. We explicitly include an external Zeeman field $B$ along the $\hat{\bm{z}}-$direction in Eq. (II) and for simplicity we define $h\equiv g\mu_{B}B$, where $g$ is the $g$ factor and $\mu_{B}$ is the Bohr magneton. In contrast to Ref. [Lu2021], we have incorporated the magnetic anisotropy energy, which plays a crucial role in establishing magnetic ordering in 2D magnets, because its presence opens up an excitation gap [Gong2017].
@@
As mentioned earlier, the spectrum is gapped at Dirac points ($K_{\pm}$) due to the presence of DMI at the non-interacting level. Specifically, $h_{x}^{2}(K_{\pm})+h_{y}^{2}(K_{\pm})$ vanishes, while $h_{z}(K_{\pm})$ is nonzero. In Ref. [Owerre16], it is also found that the magnitude of the thermal Hall conductivity increases monotonically with the temperature for the non-interacting theory. However, the interaction between magnons becomes more significant when the system is no longer in the low-temperature regime, and the interaction naturally arises from the expansion of the HP transformation. To illustrate this, we first express the expansion as follows:
@@
A crucial distinction is that the operators are expanded to the order of $S^{-\frac{3}{2}}$ in the current manuscript while Ref. [Lu2021] only considers terms up to the order of $S^{-\frac{1}{2}}$. With the use of Eqs. (3), the Heisenberg model, Eq. (II), after the HP transformation, now contains both four-operator and six-operator types of interactions, in addition to the non-interacting part. The quartic interaction also contains contributions that originate from the normal-ordering of the six-operator terms; see Appendix A for details. These terms play a role in the comparison between $T_{c}$ and $T_{\mathrm{Curie}}$. For clarity, the model Hamiltonian is expressed as $\mathcal{H}\approx\mathcal{H}_{0}+\mathcal{H}_{4}+\mathcal{H}_{6}$, where $\mathcal{H}_{0}$ is defined above and $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$ denote the four-operator and six-operator interactions, respectively. The lengthy real-space and momentum-space expressions for $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$, together with their explicit derivations and mean-field decoupled forms, are given in Appendix A. In the main text we focus on the physical content of the self-consistent theory, while the Appendix retains the technical details for completeness and reproducibility.
@@
We note that if the Dyson–Maleev representation were used instead of the Holstein–Primakoff mapping, the Hamiltonian would contain only quartic interaction terms without explicit sextic contribution [Liu92]. In that formulation, the effects attributed here to $H_{6}$ would be implicitly generated by renormalizations from quartic interactions when treated self-consistently. We therefore expect the same physical renormalizations to arise in both representations, although the Holstein–Primakoff language makes the role of $H_{6}$ in driving topological transitions more transparent.
@@
We are now in a position to discuss how to solve this problem self-consistently. The main types of thermal average that must be determined in a self-consistent manner are $\langle\hat{a}^{\dagger}_{\bf q}\hat{a}_{\bf q}\rangle$, $\langle\hat{a}^{\dagger}_{\bf q}\hat{b}_{\bf q}\rangle$, and the exchange of $a$ and $b$. In doing so, the self-energy we consider can be represented by the set of diagrams shown in Fig. 1(b). In the non-self-consistent treatment, the first-order self-energy consists of a single bubble constructed from bare (non-interacting) Green’s functions. In contrast, within our self-consistent framework the same bubble is built from dressed Green’s functions, which marks a key distinction from the approach in Ref. [Lu2021]. To illustrate the role of six-operator interaction effects, Fig. 1(c) shows the second-order self-energy contribution generated by the six-operator interaction after mean-field decoupling. Note here that only normal averages are needed because the Hamiltonian preserves $U(1)$ rotation about $z$ (magnon-number conservation), which forces anomalous correlators to vanish; see Appendix A for more details.
@@
In Ref. [Pershoguba18], the sunset diagram corresponding to the second-order self-energy derived from $\mathcal{H}_{4}$ is carefully considered. The sunset diagram is not included in this work because the level of difficulty increases considerably when accounting for DMI and magnetic anisotropy. Furthermore, the self-energy associated with the sunset diagram has intrinsic complex-valued nature which further complicates the problem [Pershoguba18].
In Ref. [Lu24], the contribution from $\mathcal{H}_{6}$ is neglected, allowing for an estimate of the critical temperature $T_{c}$ based on the form of $h_{z}$, which is proportional to $m-\frac{1}{8}+\frac{m}{8S}$. Setting this expression to zero yields the condition for the topological transition, namely $m=m_{c}\equiv\frac{S}{8S+1}$. If the quartic terms with the prefactor $\frac{1}{32S^{2}}$ in Eq. (A)—arising from the normal-ordering of the sextic interactions of $\mathcal{H}_{6}$—are further neglected, one would then conclude that $T_{c}$ coincides with $T_{\rm Curie}$ [Li24]. This highlights the necessity of incorporating higher-order corrections in the Holstein-Primakoff expansion to obtain a quantitatively accurate description of both the thermal and topological phase transitions. It is worth mentioning that these two critical temperatures do not necessarily coincide in the strict $1/S$ expansion, because the Dirac gap is renormalized through a self-energy based on the bare Green’s function, and the thermal average of the latter does not directly correspond to the physical magnetization at this order.
@@ III Results
As mentioned in the Introduction, our main objective is to uncover the relationship between $T_{c}$ and $T_{\rm Curie}$. In the framework of SRSWT, the thermal averages $\frac{1}{N}\sum_{\bf k}\langle a^{\dagger}_{\bf k}a_{\bf k}\rangle$ can be interpreted as the difference between $S$ and the magnetization $m$, with $T_{\rm Curie}$ determined by setting $m=0$. However, a known issue of SRSWT is its prediction of first-order magnetic phase transitions, where $m$ drops discontinuously at $T_{\rm Curie}$ [Rastelli74]. We observe that this artifact appears in all of the cases presented in this work. Notably, in Sec. III.2, we demonstrate that the size of the spectral gap at the $\Gamma$ point—closely tied to the anisotropy energy and Zeeman field—is crucial for determining the behavior near $T_{\text{Curie}}$.

### v2

@@ I Introduction
Since the discovery of intrinsic ferromagnetism in two-dimensional (2D) materials, this subject has attracted extensive attention due to its potential applications in novel spintronics devices [1, 2]. In particular, magnons in these 2D Van der Waals ferromagnetic or antiferromagnetic materials play a pivotal role in transport properties such as thermal Hall effects [3, 4, 5]. Moreover, since magnons are charge-neutral, they can propagate without experiencing any joule heating dissipation [6]. Candidate materials include $\rm{Cr_{2}Ge_{2}Te_{6}}$ [1], $\rm{CrI_{3}}$ [7, 8], $\rm{CrBr_{3}}$ [9], $\rm{FePSe_{3}}$ [10], etc. Recent theoretical investigations of topological properties in 2D magnonic systems [11, 12, 13, 14], particularly under the influence of magnon-magnon interaction within honeycomb ferromagnets, also mark an important advancement in the understanding and manipulation of topological phases of Dirac magnons [15, 16, 17]. The influence of magnon-magnon interaction on honeycomb antiferromagnets has also been extensively investigated recently [18, 19, 4, 20].
@@
In 2D topological magnetic materials, the crucial ingredient in achieving nontrivial topological properties is the Dzyaloshinskii-Moriya interaction (DMI) [21, 15, 16, 22, 23, 24, 25, 26] which plays a role similar to the spin-orbit interaction in topological electronic systems [27]. Both Refs. [15] and [16] found the phenomenon of magnon-magnon interaction-induced topological phase transitions accompanied with sign reversals of the thermal Hall conductivity in honeycomb ferromagnets. These transitions can be induced by tuning the magnetic field or varying the temperature. Although the predicted phenomenon is similar in Refs. [15] and [16], the underlying setup is different in terms of the chosen directions of magnetic field and types of DMI. Alternatively, the topological phase transitions can also be controlled by chirality injection from boundaries [28], manipulation of the ground-state magnetization direction [29], or parametric magnon amplification scheme [30], etc. In addition to honeycomb lattices, we note here that topology in 2D magnetic materials is also important in kagome lattices [31, 32, 33, 34] and triangular kagome lattices [26].
@@
A common description of magnon behavior in ferromagnets primarily relies on spin-wave theory. When interactions between magnons are taken into account, intriguing phenomena such as topological phase transitions can emerge. The theory presented in Refs. [15, 16] employs the standard Green’s function approach; however, Ref. [15] only considers the case where the proper self-energy of the Hartree type is composed of bare Green’s function. This means that magnon-magnon interactions are not treated in a fully self-consistent manner in Ref. [15]. It is suggested in Ref. [35] that, to properly compare the temperatures of topological phase transitions ($T_{c}$) with Curie temperatures ($T_{\rm Curie}$), the self-consistently renormalized spin-wave theory (SRSWT) should be used, as done in Ref. [17]. Here, we refer to the Curie temperature as the point at which magnetic order disappears, regardless of whether the underlying transition is continuous or first-order, as may arise within SRSWT. The corresponding self-energy in SRSWT [36, 37, 38, 39, 40] is equivalent to summing over all bubble diagrams and depends on the dressed Green’s function. Even though the SRSWT allows extending the highest temperature to the order of $T_{\rm Curie}$, the theory still has some inherent drawbacks. For example, a key weakness is the predicted first-order ferromagnetic phase transition, whereas it is known to be second-order [38, 40]. This drawback arises from the violation of the kinematical constraints [41, 17] at high temperatures.
@@
To clarify the hierarchy of approximations, it is useful to contrast the different levels of spin-wave theory. In linear spin-wave theory (LSWT), magnon–magnon interactions are entirely neglected, leading to a systematic overestimate of $T_{\rm Curie}$ [40]. A strict $1/S$ expansion incorporates interaction terms, but evaluates self-energy diagrams using bare Green’s functions; this provides controlled corrections at intermediate temperatures. By contrast, SRSWT employs dressed propagators in the same diagrams, thereby extending the applicability of spin-wave theory up to $T_{\rm Curie}$ but at the expense of artifacts such as a spurious first-order–like transition, which do not arise in the strict $1/S$ expansion.
Despite these known limitations, self-consistent approaches offer a practical way to incorporate interaction effects beyond linear theory and strict $1/S$ expansions. This is particularly important in regimes where non-self-consistent spin-wave theories either fail or develop unphysical divergences, while self-consistent treatments remain well behaved and show good agreements with experiments. For instance, self-consistent Hartree-Fock approaches have been shown to eliminate unphysical singularities present in $1/S$ theory, yielding quantitative agreement with neutron scattering experimental data in CoNb2O6 [42] and resolving divergences in anisotropic honeycomb magnets [43]. Similarly, self-consistent nonlinear spin-wave formalisms have been widely adopted to investigate finite-temperature magnetic properties and phase stability in 2D van der Waals ferromagnets [44, 45], yielding results that align better with Quantum Monte Carlo calculations and experiments than those from standard mean-field or linear theory [45]. Self-consistent mean-field approximations have also proven essential for accurately modeling complex transport properties, such as the spin Seebeck effect in yttrium iron garnet (YIG) [46]. Furthermore, recent theoretical developments highlight the necessity of self-consistent regularization to properly treat singularities associated with topological transitions in magnon decay surfaces [47]. These successes underscore that the renormalization effects captured by SRSWT are physically motivated and essential for a correct description of interacting magnon systems.
@@
If the SRSWT is adopted for the problem, it was argued in Ref. [35] that the temperature at which the magnetization vanishes coincides with the Curie temperature. As a result, topological phase transitions would be inaccessible in transport or inelastic neutron scattering experiments. In the corresponding Reply [48], we incorporated the next-order contribution to the magnon self-energy arising from the Holstein–Primakoff (HP) expansion, and demonstrated that, even within the SRSWT framework, the topological transition occurs at a finite magnetization. Motivated by the findings in the Reply [48], the present work aims to present the calculation methods in greater detail, offer a more comprehensive analysis of the conditions under which topological phase transitions can occur—specifically, at temperatures below $T_{\rm Curie}$—and critically examine the applicability of SRSWT to Dirac magnons.
@@ II Methods
Figure 1: (a) The illustration of a honeycomb lattice, showing the three nearest-neighbor and next-nearest-neighbor vectors, labeled as $\bm{\delta}_{i}$ and $\bm{\zeta}_{i}$ ($i=1,2,3$), respectively. We set the out-of-plane direction to be along $\hat{\bm{z}}$. The blue and red sites are denoted as $A$ and $B$ sites, respectively. (b) Bubble diagrams corresponding to Hartree-type self-energy contributions in the self-consistent renormalized spin-wave theory. (c) Feynman diagram representing six-operators interaction terms.
@@
The first two terms in the Hamiltonian represent the conventional Heisenberg interactions for nearest-neighbor and next-nearest-neighbor pairs. The third term corresponds to the out-of-plane NNN Dzyaloshinskii-Moriya interaction, denoted by $\bm{D}_{ij}=\pm D\hat{\bm{z}}$, where the plus (minus) sign corresponds to clockwise (counterclockwise) hopping. We explicitly include an external Zeeman field $B$ along the $\hat{\bm{z}}-$direction in Eq. (II) and for simplicity we define $h\equiv g\mu_{B}B$, where $g$ is the $g$ factor and $\mu_{B}$ is the Bohr magneton. In contrast to Ref. [15], we have incorporated the magnetic anisotropy energy, which plays a crucial role in establishing magnetic ordering in 2D magnets, because its presence opens up an excitation gap [1].
@@
As mentioned earlier, the spectrum is gapped at Dirac points ($K_{\pm}$) due to the presence of DMI at the non-interacting level. Specifically, $h_{x}^{2}(K_{\pm})+h_{y}^{2}(K_{\pm})$ vanishes, while $h_{z}(K_{\pm})$ is nonzero. In Ref. [49], it is also found that the magnitude of the thermal Hall conductivity increases monotonically with the temperature for the non-interacting theory. However, the interaction between magnons becomes more significant when the system is no longer in the low-temperature regime, and the interaction naturally arises from the expansion of the HP transformation. To illustrate this, we first express the expansion as follows:
@@
A crucial distinction is that the operators are expanded to the order of $S^{-\frac{3}{2}}$ in the current manuscript while Ref. [15] only considers terms up to the order of $S^{-\frac{1}{2}}$. With the use of Eqs. (3), the Heisenberg model, Eq. (II), after the HP transformation, now contains both four-operator and six-operator types of interactions, in addition to the non-interacting part. The quartic interaction also contains contributions that originate from the normal-ordering of the six-operator terms; see Appendix A for details. These terms play a role in the comparison between $T_{c}$ and $T_{\mathrm{Curie}}$. For clarity, the model Hamiltonian is expressed as $\mathcal{H}\approx\mathcal{H}_{0}+\mathcal{H}_{4}+\mathcal{H}_{6}$, where $\mathcal{H}_{0}$ is defined above and $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$ denote the four-operator and six-operator interactions, respectively. The lengthy real-space and momentum-space expressions for $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$, together with their explicit derivations and mean-field decoupled forms, are given in Appendix A. In the main text we focus on the physical content of the self-consistent theory, while the Appendix retains the technical details for completeness and reproducibility.
@@
We note that if the Dyson–Maleev representation were used instead of the Holstein–Primakoff mapping, the Hamiltonian would contain only quartic interaction terms without explicit sextic contribution [50]. In that formulation, the effects attributed here to $\mathcal{H}_{6}$ would be implicitly generated by renormalizations from quartic interactions when treated self-consistently. We therefore expect the same physical renormalizations to arise in both representations, although the Holstein–Primakoff language makes the role of $\mathcal{H}_{6}$ in driving topological transitions more transparent.
@@
We are now in a position to discuss how to solve this problem self-consistently. The main types of thermal average that must be determined in a self-consistent manner are $\langle\hat{a}^{\dagger}_{\bf q}\hat{a}_{\bf q}\rangle$, $\langle\hat{a}^{\dagger}_{\bf q}\hat{b}_{\bf q}\rangle$, and the exchange of $a$ and $b$. In doing so, the self-energy we consider can be represented by the set of diagrams shown in Fig. 1(b). In the non-self-consistent treatment, the first-order self-energy consists of a single bubble constructed from bare (non-interacting) Green’s functions. In contrast, within our self-consistent framework the same bubble is built from dressed Green’s functions, which marks a key distinction from the approach in Ref. [15]. To illustrate the role of six-operator interaction effects, Fig. 1(c) shows the second-order self-energy contribution generated by the six-operator interaction after mean-field decoupling. Note here that only normal averages are needed because the Hamiltonian preserves $U(1)$ rotation about $z$ (magnon-number conservation), which forces anomalous correlators to vanish; see Appendix A for more details.
@@
In Ref. [11], the sunset diagram corresponding to the second-order self-energy derived from $\mathcal{H}_{4}$ is carefully considered. The sunset diagram is not included in this work because the level of difficulty increases considerably when accounting for DMI and magnetic anisotropy. Furthermore, the self-energy associated with the sunset diagram has an intrinsic complex-valued nature which further complicates the problem [11]. However, as detailed in the decomposition of self-energy contributions presented in Appendix C, the diagrams retained in our formalism (Fig. 1)—representing the dominant Hartree interaction channels up to order $1/S^{2}$—are sufficient to capture the essential physical competition between the softening effects of quartic terms ($H_{4}$) and the stabilizing ‘hardening’ driven by sextic corrections ($H_{6}$). Consequently, while scattering terms leading to lifetime damping are neglected, the current approximation effectively encapsulates the relevant thermodynamics required to stabilize the magnetic order.
In Ref. [48], the contribution from $\mathcal{H}_{6}$ is neglected, allowing for an estimate of the critical temperature $T_{c}$ based on the form of $h_{z}$, which is proportional to $m-\frac{1}{8}+\frac{m}{8S}$. Setting this expression to zero yields the condition for the topological transition, namely $m=m_{c}\equiv\frac{S}{8S+1}$. If the quartic terms with the prefactor $\frac{1}{32S^{2}}$ in Eq. (A)—arising from the normal-ordering of the sextic interactions of $\mathcal{H}_{6}$—are further neglected, one would then conclude that $T_{c}$ coincides with $T_{\rm Curie}$ [35]. This highlights the necessity of incorporating higher-order corrections in the Holstein-Primakoff expansion to obtain a quantitatively accurate description of both the thermal and topological phase transitions. It is worth mentioning that these two critical temperatures do not necessarily coincide in the strict $1/S$ expansion, because the Dirac gap is renormalized through a self-energy based on the bare Green’s function, and the thermal average of the latter does not directly correspond to the physical magnetization at this order.
@@ III Results
As mentioned in the Introduction, our main objective is to uncover the relationship between $T_{c}$ and $T_{\rm Curie}$. In the framework of SRSWT, the thermal averages $\frac{1}{N}\sum_{\bf k}\langle a^{\dagger}_{\bf k}a_{\bf k}\rangle$ can be interpreted as the difference between $S$ and the magnetization $m$, with $T_{\rm Curie}$ determined by setting $m=0$. However, a known issue of SRSWT is its prediction of first-order magnetic phase transitions, where $m$ drops discontinuously at $T_{\rm Curie}$ [38]. We observe that this artifact appears in all of the cases presented in this work. Notably, in Sec. III.2, we demonstrate that the size of the spectral gap at the $\Gamma$ point—closely tied to the anisotropy energy and Zeeman field—is crucial for determining the behavior near $T_{\text{Curie}}$.